# Localized necking under global compression in two-scale metallic hierarchical solids

### v1

@@ Methods
[…]
@@
Metallic hexagon-triangle hierarchical (HTH) lattices and diamond-triangle hierarchical lattices (DTH) are considered in the present study. A constituent unit cell of HTH, and the tessellation of cells that define the lattice are shown in Figs. 2(a) and 2(b) respectively. The key geometric parameters of the unit cell in Fig. 2(a) are the thickness $t$ of microscale members, the side $\ell$ of the triangular substructures, the hierarchical length-scale ratio $r$, which is the number of substructures along the length of macroscale cell wall, and $N$, the number of substructure units along its thickness. $N$ = 2 for all the simulations in this paper. The nominal slenderness $\ell/t$ of the microscale members is denoted by $\lambda$. The relative density $\rho$ of HTH is given in terms of the parameters $r$, $N$, and $\lambda$ as:
@@
Figure 6: (a) Normalized stress $\sigma_{1}^{*}/\sigma_{y}$ in a representative macro-cell wall at global strain of $\varepsilon=0.32\%$. (b & c) Zoomed-in views of the stress distribution ($\sigma_{1}^{*}/\sigma_{y}$) in the boxed, near-junction zone in panel (a) at global strains of $\varepsilon=0.32\%$ $\varepsilon=1.45\%$ respectively. (d) FE stress distribution $\sigma_{1}^{*}/\sigma_{y}$ along the s-coordinate line marked in panel (a). Here, $\rho=0.1$, $r=7$, $\sigma_{y}=181.5$ MPa.
@@
The frames (c–d) in Fig. 7 show the deformation history of members in zone ‘U’ in Fig. 7(b) as the imposed strain $\varepsilon$ increases from $0.04$ to $0.15$. Again, the superimposed colour depicts $\overline{\epsilon}$. Fig. 7(c) shows that multiple members ($c_{1},v_{1},m_{1},c_{2},v_{2},m_{2}$) have buckled plastically, unlike the NB mode in Figs. 5(a-i)–(a-iii) which involves buckling of the critical member ‘$c$’ only. In fact, the central junction ‘J’ rotates, forcing the members connected to it ($c_{1},v_{1},c_{2},v_{2}$) to buckle in a coordinated fashion. The term ‘coordinated buckling’ (CB) will thus be used to collectively refer to deformation modes that involve buckling of multiple members without any member undergoing necking. Plastic straining is confined to the ends and mid-span of these buckled members, with $\overline{\epsilon}$ attaining a maximum value of 0.48. Upon further loading to $\varepsilon=0.15$, the buckled members develop inter-member contacts and self-contacts, eventually folding upon themselves smoothly as seen in Fig. 7(d). Fig. 7(e) shows a zoomed-in view of the boxed zone ‘V’ in Fig. 7(d). A maximum $\overline{\epsilon}$ of 1.08 is attained in the concave face of the folded member ‘$v_{1}$’. The extent of plastic strain localization seen in Figs. 7(d) and (e) is remarkable; even microscale members that participate in coordinated buckling (like ‘$v_{1}$’ and ‘$c_{2}$’) have large unyielded regions. Again, the importance of continuum modelling to accurately resolve the complex pattern of contacts, and the highly nonlinear, through-thickness distribution of $\overline{\epsilon}$ in members like ‘$v_{1}$’ and ‘$c_{1}$’ in Fig. 7 will be realized. Supplementary video M3 [50] shows the evolution of deformation, plastic strain, and the stresses in members in zone ‘U’ over the entire simulation for the CB mode.
Notably, truss action prevails prior to the onset of the CB mode as well, and the CB mode is also initiated by plastic buckling of maximally stressed compression members like ‘$c_{1}$’ and ‘$c_{2}$’ in Fig. 7(c). However, the ensuing load redistribution occurs in way that prevents necking in the tension member ‘$t$’. Fig. 8(a) shows the $\sigma_{1}^{*}/\sigma_{y}$ (see Eq. (3.1)) distribution at a typical junction of an $r=7,\rho=0.06$ HTH lattice at a global strain of 1.5%. Note that at this strain, the members ‘$c$’, ‘$v$’, and ‘$m$’ are visibly buckled, and none of them, including ‘$t$’, is in a uniform state of stress.
@@
[…]
@@
The necking-buckling (NB) mode is seen in HTH lattices over a range of densities and scale ratios. For instance, Fig. 9(a) shows the deformed geometry of a low-density ($\rho=0.06,r=11$) HTH lattice at a global $\varepsilon$ of 0.14, with the superimposed colour depicting $\overline{\epsilon}$. Evidently, the NB mode remains the principal mechanism enabling rotations of the macro-cell walls about the vertices of the hexagonal macro-cells. Interestingly, as the deformation is more localized than the $\rho=0.1$ case, multiple NB pairs ($(c_{1},t_{1})$ and $(c_{2},t_{2})$ in Fig. 9(a)) are formed in the vicinity of the junctions, thus accommodating greater relative rotation between macroscale beams. The buckled member ‘$c_{1}$’ folds in upon itself forming a self-contact.
@@
[$\underset{n\times m}{[\mathbf{B}]}\,\underset{m\times 1}{\{\mathbf{F}\}}=\underset{n\times 1}{\{\mathbf{P}\}}$ (4.1)
Here $[\mathbf{B}]$ is the (rank-deficient) coefficient matrix, $\{\mathbf{P}\}$ is the vector of applied nodal forces, and $s=m-n$ is the degree of static indeterminacy. The general solution to Eq. (4.1) is given as [53, 54]
@@
where $\{\mathbf{F}_{0}\}$ is a particular solution that satisfies $[\mathbf{B}]\{\mathbf{F}_{0}\}=\{\mathbf{P}\}$, $\{\mathbf{R}\}$ is the vector of redundant forces, and the columns of $[\mathbf{B}_{1}]$ constitute a basis for the nullspace of $[\mathbf{B}]$. Physically, the $i^{th}$ column of $[\mathbf{B}_{1}]$ gives the member forces in the primary structure under the application of a unit force in the $i^{th}$ redundant. Therefore, the incompatibility $\Delta_{R}^{i}$ of the $i^{th}$ redundant member can be calculated using the unit virtual load method [54] as follows:
@@
Here $[\mathbf{G}]$ is the $m\times m$ diagonal matrix of member flexibilities. The compatibility conditions (for $i=1,2,...,s$) are therefore expressed in matrix-vector form as follows:
@@
[[$\begin{split}&\underset{m\times m}{\begin{bmatrix}[\mathbf{B}]\\ [\mathbf{B}_{1}]^{T}\,[\mathbf{G}]\end{bmatrix}}\,\underset{m\times 1}{\{\mathbf{F}\}}=\underset{m\times 1}{\begin{bmatrix}\{\mathbf{P}\}\\ \{\boldsymbol{0}\}\end{bmatrix}}\end{split}$ (4.6)
subject to the constraint Eq. (4.5). Using the elimination procedure in [53], the matrices $[\mathbf{B}_{1}]$ and $[\frac{\partial\boldsymbol{F}}{\partial\boldsymbol{P}}]$ are now computed.
@@
$02\,M_{c}(\gamma_{e})+2\,M_{v}(\gamma_{e})+M_{b}(\gamma_{e})=$ 0 (4.11)
@@ Discussion
[…]
@@ Conclusion
$\sigma_{s}$ (MPa) $\sigma_{0}$ (MPa) $N$ $\bar{\epsilon}_{0}$ $p$ $C_{1}$ $C_{2}$ $\dot{\bar{\epsilon}}_{1}$ (s-1) $\dot{\bar{\epsilon}}_{2}$ (s-1)
@@
The evolution of FE energies (including KE, as mentioned in Sec. 2.3) is useful in estimating the overall simulation quality. Table 4 presents the FE energies of a representative $r=7,\rho=0.10$ HTH lattice simulation at increasing strain levels. The energies are normalized with respect to the total external work done. Throughout the simulation, the artificial energy (AE) is maintained at a small fraction ($<10^{-2}$) of the net mechanical work done. AE is associated with hourglass control [61] of the reduced-integrated quadrilateral (CPE4R) elements in Abaqus, and its small value signifies that the mesh is well-refined. Further, the kinetic energy (KE) also remains at less than $10^{-3}$ of the work done at any point of the simulation, which indicates the absence of inertial effects despite mass-scaling [62]. Lastly, most of the external work done is dissipated by plasticity (PD) with a small fraction stored as elastic strain energy; this is characteristic of lattices in the range of densities examined in the present work.

### v2

@@ Methods
$\mathcal{E}=\sum\limits_{i}^{\begin{subarray}{c}\text{nodal}\\ \text{variables}\end{subarray}}\int\limits_{\mathcal{R}}\left(v_{i}^{\text{new}}-v_{i}^{\text{old}}\right)^{2}dA+\sum\limits_{j}^{\begin{subarray}{c}\text{IP}\\ \text{variables}\end{subarray}}\int\limits_{\mathcal{R}}\left(q_{j}^{\text{new}}-q_{j}^{\text{old}}\right)^{2}dA$ (2.1)
@@
Metallic hexagon-triangle hierarchical (HTH) lattices and diamond-triangle hierarchical lattices (DTH) are considered in the present study. A constituent unit cell of HTH, and the tessellation of cells that define the lattice are shown in Figs. 2(a) and 2(b) respectively. The key geometric parameters of the unit cell in Fig. 2(a) are the thickness $t$ of microscale members, the side $\ell$ of the triangular substructures, the hierarchical length-scale ratio $r$, which is the number of substructures along the length of macroscale cell wall, and $N$, the number of substructure units along its thickness. $N$ = 2 for all the simulations in this paper. The nominal slenderness $\ell/t$ of the microscale members is denoted by $\lambda$. The relative density $\rho$ of HTH is given in terms of the parameters $r$, NN, and $\lambda$ as:
@@
Figure 6: (a) Normalized stress $\sigma_{1}^{*}/\sigma_{y}$ in a representative macro-cell wall at global strain of $\varepsilon=0.32\%$. (b & c) Zoomed-in views of the stress distribution ($\sigma_{1}^{*}/\sigma_{y}$) in the boxed, near-junction zone in panel (a) at global strains of $\varepsilon=0.32\%$ & $\varepsilon=1.45\%$ respectively. (d) FE stress distribution $\sigma_{1}^{*}/\sigma_{y}$ along the s-coordinate line marked in panel (a). Here, $\rho=0.1$, $r=7$, $\sigma_{y}=181.5$ MPa.
@@
The frames (c–d) in Fig. 7 show the deformation history of members in zone ‘U’ in Fig. 7(b) as the imposed strain $\varepsilon$ increases from $0.04$ to $0.15$. Again, the superimposed colour depicts $\overline{\epsilon}$. Fig. 7(c) shows that multiple members ($c_{1},v_{1},m_{1},c_{2},v_{2},m_{2}$) have buckled plastically, unlike the NB mode in Figs. 5(a-i)–(a-iii) which involves buckling of the critical member ‘$c$’ only. In fact, the central junction ‘J’ rotates, forcing the members connected to it ($c_{1},v_{1},c_{2},v_{2}$) to buckle in a coordinated fashion. The term ‘coordinated buckling’ (CB) will thus be used to collectively refer to deformation modes that involve buckling of multiple members without any member undergoing necking. Plastic straining is confined to the ends and mid-span of these buckled members, with $\overline{\epsilon}$ attaining a maximum value of 0.48. Upon further loading to $\varepsilon=0.15$, the buckled members develop inter-member contacts and self-contacts, eventually folding upon themselves smoothly as seen in Fig. 7(d). Fig. 7(e) shows a zoomed-in view of the boxed zone ‘V’ in Fig. 7(d). A maximum $\overline{\epsilon}$ of 1.08 is attained in the concave face of the folded member $v_{1}$’. The extent of plastic strain localization seen in Figs. 7(d) and (e) is remarkable; even microscale members that participate in coordinated buckling (like $v_{1}$’ and $c_{2}$’) have large unyielded regions. Again, the importance of continuum modelling to accurately resolve the complex pattern of contacts, and the highly nonlinear, through-thickness distribution of $\overline{\epsilon}$ in members like $v_{1}$’ and $c_{1}$’ in Fig. 7 will be realized. Supplementary video M3 [50] shows the evolution of deformation, plastic strain, and the stresses in members in zone ‘U’ over the entire simulation for the CB mode.
Notably, truss action prevails prior to the onset of the CB mode as well, and the CB mode is also initiated by plastic buckling of maximally stressed compression members like $c_{1}$’ and $c_{2}$’ in Fig. 7(c). However, the ensuing load redistribution occurs in way that prevents necking in the tension member ‘$t$’. Fig. 8(a) shows the $\sigma_{1}^{*}/\sigma_{y}$ (see Eq. (3.1)) distribution at a typical junction of an $r=7,\rho=0.06$ HTH lattice at a global strain of 1.5%. Note that at this strain, the members ‘$c$’, ‘$v$’, and ‘$m$’ are visibly buckled, and none of them, including ‘$t$’, is in a uniform state of stress.
@@
$F_{a}=\int\limits_{S}\mathbf{E}_{a}\cdot\,{\boldsymbol{\mathrm{\sigma}}}\,\mathbf{n}\,dS$ (3.2)
@@
The necking-buckling (NB) mode is seen in HTH lattices over a range of densities and scale ratios. For instance, Fig. 9(a) shows the deformed geometry of a low-density ($\rho=0.06,r=11$) HTH lattice at a global $\varepsilon$ of 0.14, with the superimposed colour depicting $\overline{\epsilon}$. Evidently, the NB mode remains the principal mechanism enabling rotations of the macro-cell walls about the vertices of the hexagonal macro-cells. Interestingly, as the deformation is more localized than the $\rho=0.1$ case, multiple NB pairs ($(c_{1},t_{1})$ and $(c_{2},t_{2})$ in Fig. 9(a)) are formed in the vicinity of the junctions, thus accommodating greater relative rotation between macroscale beams. The buckled member $c_{1}$’ folds in upon itself forming a self-contact.
@@
$\underset{n\times m}{[\mathbf{B}]}\,\underset{m\times 1}{\{\mathbf{F}\}}=\underset{n\times 1}{\{\mathbf{P}\}}$ (4.1)
Here $\mathbf{B}]$ is the (rank-deficient) coefficient matrix, $\{\mathbf{P}\}$ is the vector of applied nodal forces, and s=mns=m-n is the degree of static indeterminacy. The general solution to Eq. (4.1) is given as [53, 54]
@@
where $\{\mathbf{F}_{0}\}$ is a particular solution that satisfies $\mathbf{B}]\{\mathbf{F}_{0}\}=\{\mathbf{P}\}$, $\{\mathbf{R}\}$ is the vector of redundant forces, and the columns of $\mathbf{B}_{1}]$ constitute a basis for the nullspace of $\mathbf{B}]$. Physically, the $i^{th}$ column of $\mathbf{B}_{1}]$ gives the member forces in the primary structure under the application of a unit force in the $i^{th}$ redundant. Therefore, the incompatibility $\Delta_{R}^{i}$ of the $i^{th}$ redundant member can be calculated using the unit virtual load method [54] as follows:
@@
Here $\mathbf{G}]$ is the $m\times m$ diagonal matrix of member flexibilities. The compatibility conditions (for i=1,2,,si=1,2,...,s) are therefore expressed in matrix-vector form as follows:
@@
$\begin{split}&\underset{m\times m}{\begin{bmatrix}[\mathbf{B}]\\ [\mathbf{B}_{1}]^{T}\,[\mathbf{G}]\end{bmatrix}}\,\underset{m\times 1}{\{\mathbf{F}\}}=\underset{m\times 1}{\begin{bmatrix}\{\mathbf{P}\}\\ \{\boldsymbol{0}\}\end{bmatrix}}\end{split}$ (4.6)
subject to the constraint Eq. (4.5). Using the elimination procedure in [53], the matrices $\mathbf{B}_{1}]$ and $\frac{\partial\boldsymbol{F}}{\partial\boldsymbol{P}}]$ are now computed.
@@
$2\,M_{c}(\gamma_{e})+2\,M_{v}(\gamma_{e})+M_{b}(\gamma_{e})=0$ (4.11)
@@ Discussion
$\tilde{F}_{m}$ $=\dfrac{1}{\varepsilon_{f}}\int\limits_{0}^{\varepsilon_{f}}\tilde{F}(\varepsilon)\,d\varepsilon$ (5.2)
$W_{p}$ $=\dfrac{1}{m}\int\limits_{0}^{t_{f}}\int\limits_{V}\boldsymbol{\mathrm{\sigma}}:\boldsymbol{\dot{\epsilon}}^{p}\,dV\,dt=\dfrac{1}{m}\int\limits_{0}^{t_{f}}\int\limits_{V}\bar{\sigma}\,\dot{\overline{\epsilon}}\,dV\,dt$ (5.3)
@@ Conclusion
$\sigma_{s}$ (MPa) $\sigma_{0}$ (MPa) NN $\bar{\epsilon}_{0}$ $p$ $C_{1}$ $C_{2}$ $\dot{\bar{\epsilon}}_{1}$ (s-1) $\dot{\bar{\epsilon}}_{2}$ (s-1)
@@
The evolution of FE energies (including KE, as mentioned in Sec. 2.3) is useful in estimating the overall simulation quality. Table 4 presents the FE energies of a representative $r=7,\rho=0.10$ HTH lattice simulation at increasing strain levels. The energies are normalized with respect to the total external work done. Throughout the simulation, the artificial energy (AE) is maintained at a small fraction ($10^{-2}$) of the net mechanical work done. AE is associated with hourglass control [61] of the reduced-integrated quadrilateral (CPE4R) elements in Abaqus, and its small value signifies that the mesh is well-refined. Further, the kinetic energy (KE) also remains at less than $10^{-3}$ of the work done at any point of the simulation, which indicates the absence of inertial effects despite mass-scaling [62]. Lastly, most of the external work done is dissipated by plasticity (PD) with a small fraction stored as elastic strain energy; this is characteristic of lattices in the range of densities examined in the present work.